# FragNet: A Graph Neural Network for Molecular Property Prediction with Four Layers of Interpretability

### v1

@@ Abstract
Molecular property prediction is a crucial step in many modern-day scientific applications including drug discovery and energy storage material design. Despite the availability of numerous machine learning models for this task, we are lacking in models that provide both high accuracies and interpretability of the predictions. We introduce the FragNet architecture, a graph neural network not only capable of achieving prediction accuracies comparable to the current state-of-the-art models, but also able to provide insight on four levels of molecular substructures. This model enables understanding of which atoms, bonds, molecular fragments, and molecular fragment connections are critical in the prediction of a given molecular property. The ability to interpret the importance of connections between fragments is of particular interest for molecules which have substructures that are not connected with regular covalent bonds. The interpretable capabilities of FragNet are key to gaining scientific insights from the model’s learned patterns between molecular structure and molecular properties.
@@
Graph Neural Networks, Deep Learning, Attention, Model Interpretability
@@ 1 Introduction
Molecular property prediction is a crucial component of material discovery and design in many modern scientific applications including drug design, energy storage material discovery, catalysis and agrochemicals. Despite the availability of a large number of machine learning models[1] for molecular property prediction, there is often a trade-off in the ability of models to provide highly accurate predictions versus their ability to provide interpretability of their predictions to enable scientific insights. In this work, we introduce a graph neural network architecture called FragNet which is not only capable of achieving prediction accuracies comparable to or exceeding the accuracies of the current state-of-the-art models but also has the ability to provide insight into four levels of molecular substructures which are the atoms, bonds, fragments and fragment connections. In other words, with this model, one can understand which atoms, bonds, molecular fragments and also which connections between molecular fragments play critical role in predicting a given molecular property. By enhancing the model with the ability to reason about the connections between fragments, we provide improved representations for molecules with substructures that are not connected with regular covalent bonds, such as salts and complexes.
@@
There are several prior works which implement deep learning models which leverage attention mechanisms which can provide values for different types of molecular substructures. The AttentiveFP [2] and MoGAT [3] models provide atom level importance values. The work by Wu et al. [4] provides fragment level attention. However, their work is limited in the range of molecular properties on which it evaluates, as it only provides results for the ESOL dataset [5] using a random train-test split configuration. In contrast, our work provides prediction results for several benchmark datasets in MoleculeNet [6] while using more challenging scaffold splitting method [7]. This provides a more comprehensive and robust understanding of how the reasoning of the molecular property prediction models is affected by the target property.
@@
In addition to demonstrating the strong property prediction performance of the FragNet model across this challenging set of tasks, we also demonstrate the utility of the multi-layered interpretability mechanisms of the through several case studies on a selection of property prediction tasks. We use model attention weights and contribution values to investigate the reasoning used by the model for both individual molecular predictions as well as aggregated across multiple predictions, to identify the key molecular constituents that drive property variations. To further validate the reasoning extracted from the models, we perform a comparative study of FragNet contribution scores with density functional theory (DFT) computations of electrostatic surface potentials. Finally, we develop and release a interactive browser application to make these types of interpretability studies accessible for other molecular property tasks.
@@ 2 Results
2.1 Model
Table 1: FragNet performance measured with RMSE on regression tasks in the MoleculeNet benchmark compared with four state-of-the-art baselines. Best results in bold. Lower is better. Reported here are mean $\pm$ standard deviation corresponding to three random seeds. Except for SimSGT, results for other models were obtained from the work of reference [8].
@@ 2.1 Model
ContextPred[9, 8] 1.196 $\pm$ 0.037 0.702 $\pm$ 0.020 1.243 $\pm$ 0.025
AttrMask[9, 8] 1.112 $\pm$ 0.048 0.730 $\pm$ 0.004 1.256 $\pm$ 0.000
JOAO[10, 8] 1.120 $\pm$ 0.019 0.708 $\pm$ 0.007 1.293 $\pm$ 0.003
GraphMVP[11, 8] 1.064 $\pm$ 0.045 0.691 $\pm$ 0.013 1.2228 $\pm$ 0.001
Mole-BERT[8] 1.015 $\pm$ 0.030 0.676 $\pm$ 0.017 1.232 $\pm$ 0.009
SimSGT[12] 0.917 $\pm$ 0.028 0.670 $\pm$ 0.015 1.036 $\pm$ 0.022
@@
ContextPred[8] 74.0 $\pm$ 3.4 59.7 $\pm$ 1.8 73.6 $\pm$ 0.3
JOAO[10, 8] 66.6 $\pm$ 3.1 60.4 $\pm$ 1.5 74.8 $\pm$ 0.6
AttrMask[9, 8] 73.5 $\pm$ 4.3 60.5 $\pm$ 0.9 75.1 $\pm$ 0.9
MGSSL[13, 8] 77.1 (4.5) 61.6 (1.0) 75.2 (0.6)
GraphMVP[11, 8] 79.1 $\pm$ 2.8 60.2 $\pm$ 1.1 74.9 $\pm$ 0.8
Mole-BERT[8] 78.9 $\pm$ 3.0 62.8 $\pm$ 1.1 76.8 $\pm$ 0.5
SimSGT[12] 85.7 $\pm$ 1.8 61.7 $\pm$ 0.8 76.8 $\pm$ 0.9
@@
The FragNet model is based on a message-passing graph neural network architecture which reasons over four different graph-based representations of the molecular structure. An overview of the graph-based representations used by our model is shown in Figure 1. The data representation used by our model consists of four graph structures: atom-based, bond-based, fragment-based and fragment connection-based. The nodes and edges of the atom graph are atoms and covalent bonds respectively. In the bond graph, the nodes are bonds, while an edge in this graph is formed when two bonds have a common atom. The fragment graph is created by decomposing the molecules into substructures using a fragmentation scheme. In this work, we used BRICS fragmentation[14, 15, 16]. If the molecule is not a salt or a complex, the edges in the fragment graph are formed by bonds at which the fragments are split.
In the case of molecular complexes, where regular bonds do not exist between some fragments, we form ‘virtual’ connections to enable message passing between such disconnected substructures. Specifically, we create virtual connections between fragments that are linked by BRICS fragments. If a molecule cannot be decomposed into fragments, we treat the molecule itself as a fragment and establish a self-connection. For molecules composed of substructures that are not connected by covalent bonds, we create virtual connections by linking each fragment in one substructure to each fragment in the other.
The FragNet model leverages a hierarchical approach among these graph-representations, using the learned representations of lower-level structures to initialize the features of the high-level structures in the each subsequent graph representation. Our model starts by creating the representation for a given molecule using the bond graph. The nodes in the bond graph are initially featurized using bond properties, while the edges in the bond graph are featurized using bond angles. As bond properties we considered bond type (single, double, triple, aromatic), whether the bond is considered to be conjugated, whether the bond is in a ring and stereochemistry of the bond. The node representation of the bond graph is updated using the graph attention mechanism [17]. The updated node features of the bond graph are then used as the initial edge features of the atom graph, as indicated by the yellow arrow in Figure 1(a). Using these edge features, the atom graph is updated, also using the graph attention mechanism. Once updated, a molecular representation is created by summing all the atom feature vectors of the atom graph.
The fragment graph is updated following a similar procedure. For the initial node features of the fragment graph, we use the summed atom features from the atom graph corresponding to the atoms of which the fragment is composed. The fragment graph node representations are then updated through the graph attention mechanism. Then the edge features of the fragment graph are initialized using the learned node features of the fragment connection graph before applying graph attention. Using the updated fragment graph a second molecular representation is constructed. The final molecular representation is the concatenation of the representations constructed based on the atom and fragment graphs. This representation can then be used for downstream tasks like molecular property prediction. A complete mathematical description of the FragNet architecture is provided in the Methods section.
2.1.1 Prediction Accuracies
We test the performance of FragNet on multiple molecular property prediction tasks derived from the MoleculeNet benchmark including four regression tasks and three classification tasks. These tasks spanned a range of chemical, biological, and toxicity properties. Before performing property prediction training, the model was pre-trained on a set of self-supervised tasks as described in the Methods section. Table 1 and Table 2 presents the prediction accuracies of FragNet obtained based on scaffold splitting[18]. For each dataset, FragNet’s hyperparameters were optimized based on the validation accuracy using three random seeds. Using the optimized parameters and the same three random seeds, FragNet was fine-tuned, and the mean and standard deviation of the prediction accuracy on the test set are reported in Tables 1 and 2. We compare the performance of FragNet with four state-of-the-art baseline approaches. With the exception of CEP and Malaria, FragNet achieves accuracies comparable to or better than the current state-of-the-art for both regression and classification tasks. For CEP, the slightly lower accuracies may be attributed to fact that we only the first hyperparameter combination suggested by Optuna compared to thirty hyperparameter optimization runs conducted for the other datasets. This is due to the time constraints of training as the training set of CEP is large, with 23,982 molecules.
@@ 2.2 Interpretability
The primary distinction between FragNet and other models lies in its ability to interpret predictions based on four different types of substructures. Notably, FragNet can handle molecules with substructures that are not connected via covalent bonds, such as salts and complexes. We can analyze all four substructure types using two different mechanisms from the model - attention weights and contribution values. Attention weights are derived from the graph attention mechanisms from each of the four graph representations and provide insight into which substructures the model focuses on when making its predictions. Meanwhile, we can quantify the contribution of a given substructure to the prediction of a molecular property value [4]. This is achieved by first predicting the property for the entire molecule in the usual manner. Subsequently, another prediction is made after masking the substructure of interest, meaning the node features of this substructure are excluded from the prediction process. If the prediction with the masked substructure is lower than that of the unmasked molecule, the masked substructure is considered property-increasing. Conversely, if the prediction with the masked substructure is higher, the masked substructure is deemed to reduce the property value. Formally, the contribution is defined as follows:
@@ 2.2.1 Interpretation Case Study
The attention weights and contribution values of 1-Naphthyloxyethylethyl-beta-chloroethylamine hydrochloride from the solubility prediction model are illustrated in Figure 2. The largest atom attention weights are on the atoms in the vicinity of positively charged nitrogen, N(2) atom. This is expected as this region likely interacts most significantly with the solvent water molecules. The next largest weights are on four atoms in the aromatic ring and ether O(8). A significant attention on ether oxygen is also understandable as it can act as a hydrogen bond acceptor allowing it to interact with water molecules. These ether oxygen can also withdraw electrons from the adjacent -C2H2- group and donate them to double ring system[19].
@@
In terms of bond attention weights, the largest weights are on the C(0)-C(1), C(3)-C(4), and C(6)-C(7) bonds. One atom of each of these bonds are connected to N(2)+. Note that these three bonds are part of fragments 0, 2, and 3. Examining the 3D structure of this molecule, we observe that these three fragments could cause steric hindrance to N(2)+. We hypothesize that the large weights on C(0)-C(1) and C(3)-C(4) could be attributed to two factors: (a) the steric effects caused by fragments 0, 2, and 3, and (b) the electron flow towards the N(2)+ atom, as alkyl groups generally exhibit electron-donating properties. It should also be noted that due to the presence of chlorine in fragment 2, its electron-donating capability might be less than that of fragment 0. Moreover, there could be an electron flow through C(6)-C(7) towards the double ring system because of the ether oxygen bonded to the fragment 3. Ethers are known to exhibit both electron-withdrawing and electron-donating effects[19]. In this molecule, it is possible that the ether oxygen is withdrawing electrons from fragment 3 and donating them to the double ring system. The large weight on the C(6)-C(7) bond could signify this interesting dynamic in electron flow, as it is situated between two electron-withdrawing atoms, O(8) and N(2).
@@
The fragments containing N+ and Cl- are the ones with the highest attention weights. This result is not surprising, as these ionic species tend to interact strongly with solvent water molecules. The strongest connection weight is observed between the N+--H fragment and the --C2H2-- group adjacent to the ether oxygen. Although the exact reasons for this interaction remain unclear, it is possible that the --C2H2-- group contributes to the stability of the compound by shielding the N+ atom from the electron-withdrawing effects of the ether oxygen. Additionally, Cl- shows significant interaction with the double ring system, while the ether oxygen interacts robustly with its adjacent hydrocarbon group. Subsequent large interaction weights are noted between the ether oxygen and Cl-, N+ and its adjacent chloroethyl group, N+ and its adjacent ethyl group as well as between N+ and Cl-.
@@
In Figure 2(d), we show the contribution values for the molecule fragments which provide information about the directional impact of different substructures on the predicted property. From the fragment contributions, we identify that the double ring system is the most hydrophobic based on its highly negative contribution value indicating that it significantly decreases the model’s predicted solubility. The next most hydrophobic groups are fragments 2 and 0. The hydrophobicity of these two fragments can be mainly attributed to the steric effects they impose on N+. Given their ability to form hydrogen bonds with solvent water, it is not surprising that the most hydrophilic substructures are Cl- and those containing oxygen and N+.

### v2

# FragNet: A Graph Neural Network for Molecular Property Prediction with Four Levels of Interpretability

@@ Abstract
Molecular property prediction is essential in a variety of contemporary scientific fields, such as drug development and designing energy storage materials. Although there are many machine learning models available for this purpose, those that achieve high accuracy while also offering interpretability of predictions are uncommon. We present a graph neural network that not only matches the prediction accuracies of leading models but also provides insights on four levels of molecular substructures. This model helps identify which atoms, bonds, molecular fragments, and connections between fragments are significant for predicting a specific molecular property. Understanding the importance of connections between fragments is particularly valuable for molecules with substructures that do not connect through standard bonds. The model additionally can quantify the impact of specific fragments on the prediction, allowing the identification of fragments that may improve or degrade a property value. These interpretable features are essential for deriving scientific insights from the model’s learned relationships between molecular structures and properties.
@@
American Chemical Society,
[…]
@@ 1 Introduction
Molecular property prediction is a pivotal aspect of material discovery and design in a wide range of modern scientific fields, including drug design, energy storage material discovery, catalysis, and agrochemicals. Despite the availability of numerous machine learning models 1, 2, 3, there is often a trade-off between achieving highly accurate predictions and ensuring the interpretability of these predictions, which is crucial for scientific insights. The ability to interpret machine learning models is particularly important in the context of molecular property prediction because it allows scientists to understand the underlying factors that drive these predictions 4, 5. This understanding can lead to the identification of new hypotheses, the discovery of novel materials with desired properties, and the acceleration of the molecular design process. Furthermore, interpretability is also helpful to understand the strengths and limitations of models and thereby develop improvements 6.
Owing to these benefits, a surge of interest has sparked in explainable Artificial Intelligence (XAI), which aims to provide a deeper understanding of the model’s prediction process 7, 8. The terms explainability and interpretability have been used interchangeably in the literature to define explainable AI. Biran and Cotton 9 define the interpretability of a model as “the degree to which an observer can understand the cause of a decision.” Nigam Shah 10 categorizes interpretability into three types: engineers’ interpretability, causal interpretability, and trust-inducing interpretability. Engineers’ interpretability focuses on understanding how the AI model arrived at its output or how the model functions internally. It examines the technical aspects of model architecture, algorithms, and computational processes. This approach helps developers and researchers understand data flow and transformations within the model, enabling them to debug, optimize, and refine AI systems. Causal interpretability explains why AI models make specific decisions or predictions. It identifies key factors influencing the model’s output and explores causal relationships between these factors and the results. This approach provides intuitive explanations accessible to both technical and non-technical stakeholders, helping build trust in AI systems. Trust-inducing interpretability is concerned with fostering confidence in the model’s predictions. This approach aims to present the model’s decision-making process in a way that is understandable and relatable to users, regardless of their technical expertise. It involves creating clear, accessible explanations of how the model arrives at its conclusions, often using simplified language or visual representations. In essence, an explainable or interpretable model is one that can provide the user with some kind of understanding regarding its underlying decision-making process.
Interpretable models can be categorized in several ways. One categorization identifies models as intrinsic or extrinsic (post-hoc), based on whether interpretability is integrated into the model’s prediction algorithm or not 11, 12. Extrinsic interpretability models offer the advantage of being applicable to any complex model, regardless of its architecture 13. They provide valuable insights into the behavior of black-box models without the need to alter the models themselves, making them particularly useful for validating complex algorithms and increasing trust in their outputs 14. Interpretation methods such as SHAP (SHapley Additive exPlanations) 12 and LIME (Local Interpretable Model-agnostic Explanations) 15 are widely used for explaining traditional machine learning models by attributing feature importance scores. SHAP uses Shapley values from cooperative game theory to compute the contribution of each feature, while LIME approximates the model’s behavior around a specific prediction with an interpretable surrogate model. While powerful for typical tabular and image data, these methods face challenges when applied to graph-based molecular data. Molecules represented as graphs contain interconnected atoms and bonds, making it difficult for SHAP and LIME to capture hierarchical relationships and specialized interactions unique to molecular properties.
Intrinsic interpretability is often more advantageous in domains where transparency and simplicity are paramount. Intrinsic models are inherently easier to understand and explain. Their transparent decision-making processes make them particularly appealing in critical areas such as healthcare and finance, where the ability to directly interpret and trust the model’s decisions is essential 8. Intrinsic interpretable models have been developed for models that use different molecular representations 16. These representations include SMILES strings, molecular fingerprints, and descriptors, each offering different levels of granularity and interpretability. Feature importance values or attention weights on SMILES tokens either do not provide a comprehensive understanding of the importance of specific structural elements of the molecule or are difficult to interpret. Additionally, creating a unified interpretability functionality for different molecular representations (such as SMILES encoding, molecular descriptors, etc.) is challenging. Molecules are inherently graph-like structures, with atoms as nodes and bonds as edges. This alignment makes Graph Neural Networks (GNNs) a natural representation for molecules, allowing them to effectively capture the spatial relationships and connectivity patterns within molecules 3, 17.
@@
Previous work on interpretable GNNs for molecular property prediction has led to exciting developments, such as models designed to highlight important substructures or provide attention maps indicating which parts of the molecule contribute most to the prediction 18, 19. The AttentiveFP 20 and MoGAT 21 models provided atom-level importance values. The work by Wu et al. 22 provided fragment-level attention. However, their work is limited in the range of molecular properties on which it evaluates, as it only provides results for the ESOL dataset 23 using a random train-test split configuration.
Despite these advancements, there remain significant limitations. Current GNN models do not take into account the interactions between molecular substructures that are not connected by covalent bonds, such as salts and complexes. Such compounds play an important role in energy and pharmaceutical applications. Furthermore, to the best of our knowledge, there is no interpretable GNN model that takes into account atom, bond, and molecular fragments within a single framework when providing interpretations. Most models either focus on atoms or molecular fragments 24.
In this work, we introduce a novel graph neural network architecture named FragNet. FragNet not only achieves prediction accuracies comparable to those of state-of-the-art models, but it also provides insightful analysis across four levels of molecular substructures: atoms, bonds, fragments, and fragment connections. Here, a ‘fragment’ refers to a distinct substructure within a molecule, comprising groups of atoms and the covalent bonds that connect them. FragNet’s architecture includes graphs for each type of substructure, thereby enabling an understanding of which atoms, bonds, molecular fragments, and connections between fragments are critical in predicting a given molecular property. By constructing a fragment-based graph, we facilitate message passing between fragments that are not directly connected by covalent bonds. To the best of our knowledge, FragNet is the first model capable of simultaneously attending to all these substructures within a single framework.
We have also implemented a method to quantify the contribution of a molecular fragment to the prediction. We show that our results on the attention values and contribution values align with the well established chemical observations, thus improving the trust of the model’s prediction process. By enhancing the model with the ability to reason about the connections between fragments, we provide a more complete representation for molecules with substructures that are not connected with regular covalent bonds, such as salts and complexes.
@@
We demonstrate the utility of the multi-layered interpretability mechanisms of the model through several case studies on a selection of property prediction tasks. We use model attention weights and contribution values to investigate the reasoning used by the model for both individual molecular predictions as well as aggregated across multiple predictions, to identify the key molecular constituents that drive property variations. To further validate the reasoning extracted from the models, we perform a comparative study of FragNet contribution scores with density functional theory (DFT) computations of electrostatic surface potentials. Finally, we develop and release a interactive browser application to make these types of interpretability studies accessible for other molecular property tasks.
For readers who may not be familiar with the molecular chemistry concepts discussed in this work, we have provided a brief overview in the Supporting Information section.
2 Method
2.1 Model
The FragNet model is based on a message-passing graph neural network architecture which reasons over four different graph-based representations of the molecular structure. An overview of the graph-based representations used by our model is shown in Figure 1. The data representation used by our model consists of four graph structures: atom-based, bond-based, fragment-based and fragment connection-based. The nodes and edges of the atom graph are atoms and covalent bonds respectively. In the bond graph, the nodes are bonds, while an edge in this graph is formed when two bonds have a common atom. The fragment graph is created by decomposing the molecules into substructures using a fragmentation scheme. In this work, we used BRICS fragmentation25, 26, 27. If the molecule is not a salt or a complex, the edges in the fragment graph are formed by bonds at which the fragments are split.
In the case of molecular complexes, where regular bonds do not exist between some fragments, we form ‘virtual’ connections to enable message passing between such disconnected substructures. Specifically, we create virtual connections between fragments that are linked by BRICS fragments. These virtual connections can represent non-covalent interactions such as van der Waals forces, hydrogen bonds and electrostatic interactions. These interactions, while not as permanent as covalent bonds, are valid and significant for molecular modeling. If a molecule cannot be decomposed into fragments, we treat the molecule itself as a fragment and establish a self-connection. For molecules composed of substructures that are not connected by covalent bonds, we create virtual connections by linking each fragment in one substructure to each fragment in the other.
The FragNet model leverages a hierarchical approach among these graph-representations, using the learned representations of lower-level structures to initialize the features of the high-level structures in the each subsequent graph representation. Our model starts by creating the representation for a given molecule using the bond graph. The nodes in the bond graph are initially featurized using bond properties, while the edges in the bond graph are featurized using bond angles. As bond properties we considered bond type (single, double, triple, aromatic), whether the bond is considered to be conjugated, whether the bond is in a ring and stereochemistry of the bond.
The node representation of the bond graph is updated using the graph attention mechanism 28. The updated node features of the bond graph, $h_{i}^{bond-graph}$, are calculated as,
$h_{i}^{bond-graph}=\sum_{j\in N(i)}\alpha_{i,j}^{bond-graph}\textbf{W}h_{j,0}^{bond-graph}.$ (1)
Here, $\alpha_{i,j}^{bond-graph}$ represents the importance of node j and the edge connecting nodes $i$ and $j$ to node $i$.
where,
The updated node features of the bond graph are then used as the initial edge features of the atom graph, as indicated by the yellow arrow in Figure 1(a). Using these edge features, the atom graph is updated, also using the graph attention mechanism. The initial node features of the atom graph are atomic number, implicit valence, formal charge, number of radical electrons, hybridization, is the atom aromatic, is the atom in a ring, total number of Hydrogen atoms and chirality.
For the message passing in the atom graph, we do consider self edges (the edges connecting an atom to itself). The edge properties of the self edges are initialized as a vector of zeros having the same dimension as $h_{i}^{bond-graph}$.
Once updated, a molecular representation is created by summing all the atom feature vectors of the atom graph.
The fragment graph is updated following a similar procedure. For the initial node features of the fragment graph, we use the summed atom features from the atom graph corresponding to the atoms of which the fragment is composed.
The fragment graph’s node representations are then updated through the graph attention mechanism. Then the edge features of the fragment graph are initialized using the learned node features of the fragment connection graph before applying graph attention. Using the updated fragment graph a second molecular representation is constructed. The final molecular representation is the concatenation of the representations constructed based on the atom and fragment graphs. This representation can then be used for downstream tasks like molecular property prediction. A complete mathematical description of the FragNet architecture is provided in the Methods section.
The details on the data and the training approach are provided in the Supporting Information.
@@ 3 Results
3.1 Prediction Accuracies
We test the performance of FragNet on multiple molecular property prediction tasks derived from the MoleculeNet benchmark including four regression tasks and three classification tasks. These tasks spanned a range of chemical, biological, and toxicity properties. Before performing property prediction training, the model was pre-trained on a set of self-supervised tasks as described in the Methods section. Table 1 and Table 2 presents the prediction accuracies of FragNet obtained based on scaffold splitting29. For each dataset, FragNet’s hyperparameters were optimized based on the validation accuracy using three random seeds. Using the optimized parameters and the same three random seeds, FragNet was fine-tuned, and the mean and standard deviation of the prediction accuracy on the test set are reported in Tables 1 and 2. We compare the performance of FragNet with four state-of-the-art baseline approaches. With the exception of CEP and Malaria, FragNet achieves accuracies comparable to or better than the current state-of-the-art for both regression and classification tasks. For CEP, the slightly lower accuracies may be attributed to the fact that we only use the first hyperparameter combination suggested by Optuna compared to thirty hyperparameter optimization runs conducted for the other datasets. This is due to the time constraints of training as the training set of CEP is large, with 23,982 molecules.
Table 1: FragNet performance measured with RMSE on regression tasks in the MoleculeNet benchmark compared with four state-of-the-art baselines. Best results in bold. Lower is better. Reported here are mean $\pm$ standard deviation corresponding to three random seeds. Except for SimSGT, results for other models were obtained from the work of reference 30.
@@ 3.1 Prediction Accuracies
ContextPred31, 30 1.196 $\pm$ 0.037 0.702 $\pm$ 0.020 1.243 $\pm$ 0.025
AttrMask31, 30 1.112 $\pm$ 0.048 0.730 $\pm$ 0.004 1.256 $\pm$ 0.000
JOAO32, 30 1.120 $\pm$ 0.019 0.708 $\pm$ 0.007 1.293 $\pm$ 0.003
GraphMVP33, 30 1.064 $\pm$ 0.045 0.691 $\pm$ 0.013 1.2228 $\pm$ 0.001
Mole-BERT30 1.015 $\pm$ 0.030 0.676 $\pm$ 0.017 1.232 $\pm$ 0.009
SimSGT34 0.917 $\pm$ 0.028 0.670 $\pm$ 0.015 1.036 $\pm$ 0.022
@@
ContextPred30 74.0 $\pm$ 3.4 59.7 $\pm$ 1.8 73.6 $\pm$ 0.3
JOAO32, 30 66.6 $\pm$ 3.1 60.4 $\pm$ 1.5 74.8 $\pm$ 0.6
AttrMask31, 30 73.5 $\pm$ 4.3 60.5 $\pm$ 0.9 75.1 $\pm$ 0.9
MGSSL35, 30 77.1 $\pm$ 4.5 61.6 $\pm$ 1.0 75.2 $\pm$ 0.6
GraphMVP33, 30 79.1 $\pm$ 2.8 60.2 $\pm$ 1.1 74.9 $\pm$ 0.8
Mole-BERT30 78.9 $\pm$ 3.0 62.8 $\pm$ 1.1 76.8 $\pm$ 0.5
SimSGT34 85.7 $\pm$ 1.8 61.7 $\pm$ 0.8 76.8 $\pm$ 0.9
@@ 3.2 Interpretability
The primary distinction between FragNet and other models lies in its ability to interpret predictions based on four different types of substructures. Notably, FragNet can handle molecules with substructures that are not connected via covalent bonds, such as salts and complexes. We can analyze all four substructure types using two different mechanisms from the model - attention weights and contribution values. Attention weights are derived from the graph attention mechanisms from each of the four graph representations and provide insight into which substructures the model focuses on when making its predictions. Meanwhile, we can quantify the contribution of a given substructure to the prediction of a molecular property value 22. This is achieved by first predicting the property for the entire molecule in the usual manner. Subsequently, another prediction is made after masking the substructure of interest, meaning the node features of this substructure are excluded from the prediction process. If the prediction with the masked substructure is lower than that of the unmasked molecule, the masked substructure is considered property-increasing. Conversely, if the prediction with the masked substructure is higher, the masked substructure is deemed to reduce the property value. Formally, the contribution is defined as follows:
@@ 3.2.1 Interpretation Case Study
The attention weights and contribution values of 1-Naphthyloxyethylethyl-beta-chloroethylamine hydrochloride from the solubility prediction model are illustrated in Figure 2. The largest atom attention weights are on the atoms in the vicinity of positively charged nitrogen, N(2) atom. This is expected as this region likely interacts most significantly with the solvent water molecules. The next largest weights are on four atoms in the aromatic ring and ether O(8). A significant attention on ether oxygen is also understandable as it can act as a hydrogen bond acceptor allowing it to interact with water molecules. These ether oxygen can also withdraw electrons from the adjacent -C2H2- group and donate them to double ring system36.
@@
In terms of bond attention weights, the largest weights are on the C(0)-C(1), C(3)-C(4), and C(6)-C(7) bonds. One atom of each of these bonds are connected to N(2)+. Note that these three bonds are part of the fragments 0, 2, and 3. Examining the 3D structure of this molecule, we observe that these three fragments could cause steric hindrance to N(2)+. We hypothesize that the large weights on C(0)-C(1) and C(3)-C(4) could be attributed to two factors: (a) the steric effects caused by fragments 0, 2, and 3, and (b) the electron flow towards the N(2)+ atom, as alkyl groups generally exhibit electron-donating properties. It should also be noted that due to the presence of chlorine in fragment 2, its electron-donating capability might be less than that of fragment 0. Moreover, there could be an electron flow through C(6)-C(7) towards the double ring system because of the ether oxygen bonded to the fragment 3. Ethers are known to exhibit both electron-withdrawing and electron-donating effects36. In this molecule, it is possible that the ether oxygen is withdrawing electrons from fragment 3 and donating them to the double ring system. The large weight on the C(6)-C(7) bond could signify this interesting dynamic in electron flow, as it is situated between two electron-withdrawing atoms, O(8) and N(2).
@@
The fragments containing N+ and Cl- are the ones with the highest attention weights. This result is not surprising, as these ionic species tend to interact strongly with solvent water molecules. The strongest connection weight is observed between the N+--H fragment and the --C2H4-- group adjacent to the ether oxygen. Although the exact reasons for this interaction remain unclear, it is possible that the --C2H4-- group contributes to the stability of the compound by shielding the N+ atom from the electron-withdrawing effects of the ether oxygen. Additionally, Cl- shows significant interaction with the double ring system, while the ether oxygen interacts robustly with its adjacent hydrocarbon group. Subsequent large interaction weights are noted between the ether oxygen and Cl-, N+ and its adjacent chloroethyl group, N+ and its adjacent ethyl group as well as between N+ and Cl-.
@@
In Figure 2(d), we show the contribution values for the molecule fragments which provide information about the directional impact of different substructures on the predicted property. From the fragment contributions, we identify that the double ring system is the most hydrophobic based on its highly negative contribution value, indicating that it significantly decreases the model’s predicted solubility. The next most hydrophobic groups are fragments 2 and 0. The hydrophobicity of these two fragments can be mainly attributed to the steric effects they impose on N+. Given their ability to form hydrogen bonds with solvent water molecules, it is not surprising that the most hydrophilic substructures are Cl- and those containing oxygen and N+.